OpenStudy (anonymous):

Solve for x: (x - 1)(3/4) = 64

OpenStudy (rosedewittbukater):

Try to get x alone on one side of the equal sign. Multiply both sides by (4/3) (x-1)(3/4)(4/3) = 64(4/3) x-1 = 256/3 Add 1 to both sides x = 256/3 + 3/3 x = 259/3